OpenStudy (anonymous):

An airplane leaves an airport and flies due west 150 miles and then 170 miles in the direction S 49.17°W. How far is the plane from the airport (round to the nearest mile)?

OpenStudy (anonymous):

Let x= distance from airport to plane in miles x^2=170^2+150^2-2(170)(150)cos(49.17+90) x^2=89,989.29 x=300 miles ^is this right?